Assignment Vocou Colculate E field eserywhore. Concent Spheres

Help me to solve it. (Physics II)
Answer #1

The problem can be solved by using Gauss's Law,

\int _S \vec{E}.d\vec{a} = \frac{q_{en}}{\epsilon _0}

where q_{en} = the enclosed charge.

Since a sphere is symmetric,

\vec{E}.d\vec{a} = Edr

And, thus,

\int _S Eda= \frac{q_{en}}{\epsilon _0}

E\int _Sda= \frac{q_{en}}{\epsilon _0}

E\left ( 4 \pi r^2 \right )= \frac{q_{en}}{\epsilon _0}

E= \frac{1}{4 \pi \epsilon _0}\frac{q_{en}}{r^2}

where r is the distance of the point of interest from center.

For position a:

q_{en} = +q

E_a= \frac{1}{4 \pi \epsilon _0}\frac{q}{r^2}

For position b:

q_{en} = +q -3q = -2q

E_b= \frac{1}{4 \pi \epsilon _0}\frac{-2q}{r^2}

For position c:

q_{en} = +q - 3q +5q = 3q

E_c= \frac{1}{4 \pi \epsilon _0}\frac{3q}{r^2}